OpenStudy (anonymous):

Simplify √-100 Answers: -10 -10 i 10 i 10

OpenStudy (angela210793):

=i*sqr100=i*10=10i

OpenStudy (anonymous):

10i

OpenStudy (anonymous):

10i

OpenStudy (angela210793):

keep goin with medals ppl :P